OpenStudy (anonymous):

For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b? (2b, 2b + 7) (2b, x + 7) (x, x + 7) (x, 2b + 7)

myininaya (myininaya):

f(2b)=?

myininaya (myininaya):

Replace everything with x in f(x)=x+7 with 2b

myininaya (myininaya):

your ordered pair will be in the form (x,f(x))

OpenStudy (anonymous):

Oh hey there, again! :D Okay erm... f(x) = x + 7 f(2b) = 2b + 7 I replaced the 'x's. :))

myininaya (myininaya):

So which do you think he is proper choice above? You already replaced x with 2b (so x=2b) And the output you got is f(2b). So just replace those in here: (x,f(x)) (2b,f(2b)) You put f(2b) in. :)